In December 2016, Learer Company’s manager estimated next year’s total direct labor cost assuming 35 persons working an average of 3,000 hours each at an average wage rate of $20 per hour. The manager also estimated the following manufacturing overhead costs for 2017.

Indirect labor $ 335,200
Factory supervision 165,000
Rent on factory building 156,000
Factory utilities 104,000
Factory insurance expired 84,000
Depreciation—Factory equipment 413,000
Repairs expense—Factory equipment 76,000
Factory supplies used 84,800
Miscellaneous production costs 52,000
Total estimated overhead costs $ 1,470,000


At the end of 2017, records show the company incurred $1,843,000 of actual overhead costs. It completed and sold five jobs with the following direct labor costs: Job 201, $620,000; Job 202, $579,000; Job 203, $314,000; Job 204, $732,000; and Job 205, $330,000. In addition, Job 206 is in process at the end of 2017 and had been charged $33,000 for direct labor. No jobs were in process at the end of 2016. The company’s predetermined overhead rate is based on direct labor cost.

Required
1-a.
Determine the predetermined overhead rate for 2017.
1-b. Determine the total overhead cost applied to each of the six jobs during 2017.
1-c. Determine the over- or underapplied overhead at year-end 2017.
2. Assuming that any over- or underapplied overhead is not material, prepare the adjusting entry to allocate any over- or underapplied overhead to Cost of Goods Sold at the end of 2017.
Answer #1

1a) Estimated direct labor cost = 35*3000*20 = 2100000

Overhead rate = 1470000/2100000 = 70% of direct labor cost

1b) Total Direct labor cost = 620000+579000+314000+732000+330000+33000 = 2608000

Overhead applied = 2608000*70% = 1825600

1c) Under applied overhead = 1843000-1825600 = 17400

2) Adjusting entry

Date account and explanation Debit credit
Cost of goods sold 17400
Manufacturing overhead 17400
(To record under applied overhead)